OpenStudy (anonymous):

What is the simplified form of 3/(g^-2)(h^3)

OpenStudy (sasogeek):

any time you have a negative exponent, it means that the base of the exponent is on the opposite side of the fraction with a positive exponent. so \(\huge \frac{3}{2^{-1}}=3*2=6 \) can you apply the same thing to your question?